A valószínűségszámítás II. gyakorlaton való részvétel

dátum:  IX. 11. IX. 18. IX. 25. X. 2. X. 9. X. 16. X. 30.

diákok:

Ádám Zsolt  +  −  −  −  −  −  −
Aradi Bernadette  +  −  +  +  +  +  +
Balázs Éva  −  −  +  −  −  +  −
Barna Timea  −  +  +  +  +  +  +
Bócsi Judit  +  +  +  +  −  +  −
Bodnár Nelli  +  +  +  +  +  +  +
Dobány Máté  +  +  +  +  +  +  +
Fista Norbert  −  −  +  +  +  +  +
Flaskár Ibolya  −  −  +  −  −  −  −
Jenei Tamás  +  +  +  +  +  +  +
Juhász Mónika  +  −  +  −  −  +  +
Karakas Éva  +  +  +  −  +  +  +
Kész Timea  +  +  +  +  +  +  −
Kontor Ibolya  −  −  +  +  +  +  +
Kovács György  −  −  +  −  +  −  +
Kovács Viktor  −  −  −  +  +  +  +
Kövér Ildikó  +  +  +  +  +  +  +
Léstár Adrienn  +  −  +  +  +  +  +
Lippai Zsolt  +  +  +  +  −  +  −
Mályi Nikoletta  +  −  +  +  +  −  +
Mezei Anita  +  +  +  +  +  +  +
Mizsei Zsolt  −  −  −  +  +  +  +
Nagy Beáta  +  +  +  −  +  +  +
Péter Antal  +  +  −  +  +  +  +
Pósa Csenge  +  +  +  +  −  +  −
Sikolya Kinga  +  −  +  +  +  +  +
Szokol Patricia  +  −  +  +  +  +  −
Szoó Tibor  +  +  +  +  −  +  −
Tóth Andrea  +  −  +  +  +  +  +
Tóth Eszter  −  −  −  −  −  −  +
Tóth Judit  −  −  +  −  −  +  −
Tóth Veronika  +  +  +  −  +  −  +
Tóth Zoltán  +  +  +  +  +  +  +
Weinémer Andrea  +  +  +  +  +  +  −

dátum:  XI. 06. XI. 13. XI. 20. XI. 27. XII. 4.

diákok:

Ádám Zsolt  −  −  −  −
Aradi Bernadette  +  +  +  +
Balázs Éva  +  +  −  −
Barna Timea  +  +  −  +
Bócsi Judit  +  +  +  −
Bodnár Nelli  +  +  +  +
Dobány Máté  +  +  +  +
Fista Norbert  +  +  +  +
Flaskár Ibolya  −  −  −  −
Jenei Tamás  +  +  +  −
Juhász Mónika  +  +  +  +
Karakas Éva  +  +  +  +
Kész Timea  +  +  +  +
Kontor Ibolya  −  +  +  +
Kovács György  +  −  +  −
Kovács Viktor  +  −  +  −
Kövér Ildikó  +  +  +  −
Léstár Adrienn  +  −  +  −
Lippai Zsolt  +  +  +  +
Mályi Nikoletta  +  −  +  −
Mezei Anita  +  +  +  +
Mizsei Zsolt  +  +  +  −
Nagy Beáta  −  −  +  +
Péter Antal  −  +  +  −
Pósa Csenge  +  +  +  −
Sikolya Kinga  +  +  +  +
Szokol Patricia  +  +  +  +
Szoó Tibor  +  +  +  −
Tóth Andrea  +  −  +  −
Tóth Eszter  +  +  +  −
Tóth Judit  −  +  −  −
Tóth Veronika  −  −  +  +
Tóth Zoltán  +  +  +  +
Weinémer Andrea  +  +  +  +